OpenStudy (anonymous):

Can someone help me set up this problem? One pump can drain a pool in 10 minutes. When a second pump is also used, the pool only takes 7 minutes to drain. How long would it take the second pump to drain the pool if it were the only pump in use?

OpenStudy (anonymous):

I think it's set up like: 1/10=x/7

OpenStudy (anonymous):

I'm not 100% sure though. But I mean, it sounds like it's set up like that.

OpenStudy (anonymous):

The time for the second pump isn't 7 though. 7 minutes is the time of both pumps working together, I think there is tad more to it...

OpenStudy (anonymous):

Oh X( I knew I was missing something.

OpenStudy (anonymous):

Thanks anyhow!

OpenStudy (anonymous):

:) You're welcome

OpenStudy (anonymous):

The time the first pump alone takes to drain the pool = 10 minutes. So it takes one minute to drain 1/10 of the pool. Let the second pump take x minutes. Then per minute capacity of draining by the second pump = 1/x of the pool. Therefore, the joint capacity of the two pumps in one minute = 1/10 + 1/x. So the per minute draining of both pumps together is 1/10 + 1/x = 1/7. Solving this equation, 1/x = 1/7 - 1/10 = (10 - 7)/70 = 3/70 Then it's a fraction equal to a fraction of 1/x = 3/70 3x = 70 and you get x = 70/3.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Sorry to solve it completely for you. I believe that's the right answer. I tried to explain as much as I could. I hope it's something similar to what you have in your notes :)